\begin{thebibliography}{33}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Anderson et~al.(1978)Anderson, Box, and Jenkins]{ARMIA}
Oliver~D. Anderson, George E.~P. Box, and Gwilym~M. Jenkins.
\newblock Time series analysis: Forecasting and control.
\newblock \emph{The Statistician}, pp.\  265, Sep 1978.
\newblock \doi{10.2307/2988198}.
\newblock URL \url{http://dx.doi.org/10.2307/2988198}.

\bibitem[Bai et~al.(2020)Bai, Yao, Li, Wang, and Wang]{AGCRN}
Lei Bai, Lina Yao, Can Li, Xianzhi Wang, and Can Wang.
\newblock Adaptive graph convolutional recurrent network for traffic forecasting.
\newblock \emph{arXiv preprint arXiv:2007.02842}, Jul 2020.

\bibitem[Bai et~al.(2018)Bai, Kolter, and Koltun]{tcn}
Shaojie Bai, J.~Zico Kolter, and Vladlen Koltun.
\newblock An empirical evaluation of generic convolutional and recurrent networks for sequence modeling.
\newblock \emph{arXiv preprint arXiv:1803.01271}, 2018.
\newblock URL \url{https://arxiv.org/abs/1803.01271}.

\bibitem[Cao et~al.(2025)Cao, Wang, Jiang, Yu, and Dong]{STDN}
Lingxiao Cao, Bin Wang, Guiyuan Jiang, Yanwei Yu, and Junyu Dong.
\newblock Spatiotemporal-aware trend-seasonality decomposition network for traffic flow forecasting, 2025.
\newblock URL \url{https://arxiv.org/abs/2502.12213}.

\bibitem[Chao et~al.(2020)Chao, Lin, Guo, and Wan]{pems}
Song Chao, Youfang Lin, Shengnan Guo, and Huaiyu Wan.
\newblock Spatial-temporal synchronous graph convolutional networks: A new framework for spatial-temporal network data forecasting.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pp.\  914–921, Jun 2020.
\newblock \doi{10.1609/aaai.v34i01.5438}.
\newblock URL \url{http://dx.doi.org/10.1609/aaai.v34i01.5438}.

\bibitem[Chung et~al.(2014)Chung, Gulcehre, Cho, and Bengio]{gru}
Junyoung Chung, Caglar Gulcehre, KyungHyun Cho, and Yoshua Bengio.
\newblock Empirical evaluation of gated recurrent neural networks on sequence modeling.
\newblock \emph{arXiv preprint arXiv:1412.3555}, 2014.
\newblock URL \url{https://arxiv.org/abs/1412.3555}.

\bibitem[Cini et~al.(2024)Cini, Zambon, and Alippi]{SGL}
Andrea Cini, Daniele Zambon, and Cesare Alippi.
\newblock Sparse graph learning from spatiotemporal time series.
\newblock \emph{Journal of Machine Learning Research}, 24\penalty0 (1), March 2024.
\newblock ISSN 1532-4435.

\bibitem[Diao et~al.(2019)Diao, Wang, Zhang, Liu, Xie, and He]{dgcnn}
Zulong Diao, Xin Wang, Dafang Zhang, Yingru Liu, Kun Xie, and Shaoyao He.
\newblock Dynamic spatial-temporal graph convolutional neural networks for traffic forecasting.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pp.\  890–897, Sep 2019.
\newblock \doi{10.1609/aaai.v33i01.3301890}.
\newblock URL \url{http://dx.doi.org/10.1609/aaai.v33i01.3301890}.

\bibitem[Dong et~al.(2024)Dong, Jiang, Gao, Liu, Deng, Wen, and Song]{HimNet}
Zheng Dong, Renhe Jiang, Haotian Gao, Hangchen Liu, Jinliang Deng, Qingsong Wen, and Xuan Song.
\newblock Heterogeneity-informed meta-parameter learning for spatiotemporal time series forecasting.
\newblock \emph{arXiv preprint arXiv:2405.10800}, 2024.
\newblock URL \url{https://arxiv.org/abs/2405.10800}.

\bibitem[Dosovitskiy et~al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]{vit}
Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, Jakob Uszkoreit, and Neil Houlsby.
\newblock An image is worth 16x16 words: Transformers for image recognition at scale.
\newblock \emph{arXiv preprint arXiv:2010.11929}, 2021.
\newblock URL \url{https://arxiv.org/abs/2010.11929}.

\bibitem[Fang et~al.(2023)Fang, Qin, Luo, Zhao, Xu, Zeng, and Wang]{STWave}
Yuchen Fang, Yanjun Qin, Haiyong Luo, Fang Zhao, Bingbing Xu, Liang Zeng, and Chenxing Wang.
\newblock When spatio-temporal meet wavelets: Disentangled traffic forecasting via efficient spectral graph attention networks.
\newblock In \emph{2023 IEEE 39th International Conference on Data Engineering (ICDE)}, pp.\  517--529. IEEE, 2023.

\bibitem[Gao et~al.(2024)Gao, Jiang, Dong, Deng, Ma, and Song]{STD-MAE}
Haotian Gao, Renhe Jiang, Zheng Dong, Jinliang Deng, Yuxin Ma, and Xuan Song.
\newblock Spatial-temporal-decoupled masked pre-training for spatiotemporal forecasting.
\newblock \emph{arXiv preprint arXiv:2312.00516}, 2024.
\newblock URL \url{https://arxiv.org/abs/2312.00516}.

\bibitem[Han et~al.(2024)Han, Zhang, Liu, Tao, Tan, and Xiong]{BigST}
Jindong Han, Weijia Zhang, Hao Liu, Tao Tao, Naiqiang Tan, and Hui Xiong.
\newblock Bigst: Linear complexity spatio-temporal graph neural network for traffic forecasting on large-scale road networks.
\newblock \emph{Proc. VLDB Endow.}, 17\penalty0 (5):\penalty0 1081–1090, January 2024.
\newblock ISSN 2150-8097.
\newblock \doi{10.14778/3641204.3641217}.
\newblock URL \url{https://doi.org/10.14778/3641204.3641217}.

\bibitem[Hochreiter \& Schmidhuber(1997)Hochreiter and Schmidhuber]{lstm}
Sepp Hochreiter and Jürgen Schmidhuber.
\newblock Long short-term memory.
\newblock \emph{Neural Computation}, 9\penalty0 (8):\penalty0 1735--1780, 1997.
\newblock \doi{10.1162/neco.1997.9.8.1735}.

\bibitem[Jiang et~al.(2023)Jiang, Han, Zhao, and Wang]{PDformer}
Jiawei Jiang, Chengkai Han, Wayne~Xin Zhao, and Jingyuan Wang.
\newblock Pdformer: Propagation delay-aware dynamic long-range transformer for traffic flow prediction.
\newblock In \emph{{AAAI}}. {AAAI} Press, 2023.

\bibitem[Kipf \& Welling(2017)Kipf and Welling]{gcn}
Thomas~N. Kipf and Max Welling.
\newblock Semi-supervised classification with graph convolutional networks.
\newblock \emph{arXiv preprint arXiv:1609.02907}, 2017.
\newblock URL \url{https://arxiv.org/abs/1609.02907}.

\bibitem[Li et~al.(2023)Li, Feng, Yan, Jin, Yang, Sun, Jin, and Li]{DGCRN}
Fuxian Li, Jie Feng, Huan Yan, Guangyin Jin, Fan Yang, Funing Sun, Depeng Jin, and Yong Li.
\newblock Dynamic graph convolutional recurrent network for traffic prediction: Benchmark and solution.
\newblock \emph{ACM Transactions on Knowledge Discovery from Data}, pp.\  1–21, Feb 2023.
\newblock \doi{10.1145/3532611}.
\newblock URL \url{http://dx.doi.org/10.1145/3532611}.

\bibitem[Li et~al.(2018)Li, Yu, Shahabi, and Liu]{DCRNN}
Yaguang Li, Rose Yu, Cyrus Shahabi, and Yan Liu.
\newblock Diffusion convolutional recurrent neural network: Data-driven traffic forecasting.
\newblock In \emph{International Conference on Learning Representations (ICLR '18)}, 2018.

\bibitem[Li et~al.(2024)Li, Shao, Xu, Qiu, Cao, and Wang]{DFDGCN}
Yujie Li, Zezhi Shao, Yongjun Xu, Qiang Qiu, Zhaogang Cao, and Fei Wang.
\newblock Dynamic frequency domain graph convolutional network for traffic forecasting.
\newblock In \emph{ICASSP 2024-2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP)}, pp.\  5245--5249. IEEE, 2024.

\bibitem[Liu et~al.(2023)Liu, Dong, Renhe, Deng, Deng, Chen, and Song]{STAEformer}
Hangchen Liu, Zheng Dong, Jiang Renhe, Jiewen Deng, Jinliang Deng, Quanjun Chen, and Xuan Song.
\newblock Spatio-temporal adaptive embedding makes vanilla transformer sota for traffic forecasting.
\newblock In \emph{Proceedings of the 32nd ACM International Conference on Information and Knowledge Management}, pp.\  4125--4129, 2023.

\bibitem[Oreshkin et~al.(2019)Oreshkin, Carpov, Chapados, and Bengio]{nbeats}
BorisN. Oreshkin, Dmitri Carpov, Nicolas Chapados, and Yoshua Bengio.
\newblock N-beats: Neural basis expansion analysis for interpretable time series forecasting.
\newblock \emph{Learning,Learning}, May 2019.

\bibitem[Shang et~al.(2021)Shang, Chen, and Bi]{GTS}
Chao Shang, Jie Chen, and Jinbo Bi.
\newblock Discrete graph structure learning for forecasting multiple time series.
\newblock \emph{arXiv preprint arXiv:2101.06861}, 2021.

\bibitem[Shao et~al.(2022{\natexlab{a}})Shao, Zhang, Wang, and Xu]{STEP}
Zezhi Shao, Zhao Zhang, Fei Wang, and Yongjun Xu.
\newblock Pre-training enhanced spatial-temporal graph neural network for multivariate time series forecasting.
\newblock In \emph{Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining}, KDD ’22, pp.\  1567–1577. ACM, August 2022{\natexlab{a}}.
\newblock \doi{10.1145/3534678.3539396}.
\newblock URL \url{http://dx.doi.org/10.1145/3534678.3539396}.

\bibitem[Shao et~al.(2022{\natexlab{b}})Shao, Zhang, Wei, Wang, Xu, Cao, and Jensen]{D2STGNN}
Zezhi Shao, Zhao Zhang, Wei Wei, Fei Wang, Yongjun Xu, Xin Cao, and Christian~S. Jensen.
\newblock Decoupled dynamic spatial-temporal graph neural network for traffic forecasting.
\newblock \emph{Proc. VLDB Endow.}, 15\penalty0 (11):\penalty0 2733–2746, July 2022{\natexlab{b}}.
\newblock ISSN 2150-8097.
\newblock \doi{10.14778/3551793.3551827}.
\newblock URL \url{https://doi.org/10.14778/3551793.3551827}.

\bibitem[Shao et~al.(2025)Shao, Wang, Xu, Wei, Yu, Zhang, Yao, Sun, Jin, Cao, Cong, Jensen, and Cheng]{exploring}
Zezhi Shao, Fei Wang, Yongjun Xu, Wei Wei, Chengqing Yu, Zhao Zhang, Di~Yao, Tao Sun, Guangyin Jin, Xin Cao, Gao Cong, Christian~S. Jensen, and Xueqi Cheng.
\newblock Exploring progress in multivariate time series forecasting: Comprehensive benchmarking and heterogeneity analysis.
\newblock \emph{IEEE Trans. on Knowl. and Data Eng.}, 37\penalty0 (1):\penalty0 291–305, January 2025.
\newblock ISSN 1041-4347.
\newblock URL \url{https://doi.org/10.1109/TKDE.2024.3484454}.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin]{attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, AidanN. Gomez, Lukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock \emph{Neural Information Processing Systems,Neural Information Processing Systems}, Jun 2017.

\bibitem[Wang et~al.(2020)Wang, Li, Khabsa, Fang, and Ma]{linformer}
Sinong Wang, Belinda~Z. Li, Madian Khabsa, Han Fang, and Hao Ma.
\newblock Linformer: Self-attention with linear complexity.
\newblock \emph{arXiv preprint arXiv:2006.04768}, 2020.

\bibitem[Wu et~al.(2019)Wu, Pan, Long, Jiang, and Zhang]{GWNet}
Zonghan Wu, Shirui Pan, Guodong Long, Jing Jiang, and Chengqi Zhang.
\newblock Graph wavenet for deep spatial-temporal graph modeling.
\newblock In \emph{Proceedings of the Twenty-Eighth International Joint Conference on Artificial Intelligence}, Aug 2019.
\newblock \doi{10.24963/ijcai.2019/264}.
\newblock URL \url{http://dx.doi.org/10.24963/ijcai.2019/264}.

\bibitem[Wu et~al.(2020)Wu, Pan, Long, Jiang, Chang, and Zhang]{MTGNN}
Zonghan Wu, Shirui Pan, Guodong Long, Jing Jiang, Xiaojun Chang, and Chengqi Zhang.
\newblock Connecting the dots: Multivariate time series forecasting with graph neural networks.
\newblock In \emph{Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining}, Aug 2020.
\newblock \doi{10.1145/3394486.3403118}.
\newblock URL \url{http://dx.doi.org/10.1145/3394486.3403118}.

\bibitem[Yu et~al.(2018)Yu, Yin, and Zhu]{STGCN}
Bing Yu, Haoteng Yin, and Zhanxing Zhu.
\newblock Spatio-temporal graph convolutional networks: A deep learning framework for traffic forecasting.
\newblock In \emph{Proceedings of the Twenty-Seventh International Joint Conference on Artificial Intelligence}, Jul 2018.
\newblock \doi{10.24963/ijcai.2018/505}.
\newblock URL \url{http://dx.doi.org/10.24963/ijcai.2018/505}.

\bibitem[Zhang \& Yan(2023)Zhang and Yan]{Crossformer}
Yunhao Zhang and Junchi Yan.
\newblock Crossformer: Transformer utilizing cross-dimension dependency for multivariate time series forecasting.
\newblock In \emph{International Conference on Learning Representations}, 2023.
\newblock URL \url{https://api.semanticscholar.org/CorpusID:259298223}.

\bibitem[Zhou et~al.(2021)Zhou, Zhang, Peng, Zhang, Li, Xiong, and Zhang]{informer}
Haoyi Zhou, Shanghang Zhang, Jieqi Peng, Shuai Zhang, Jianxin Li, Hui Xiong, and Wancai Zhang.
\newblock Informer: Beyond efficient transformer for long sequence time-series forecasting.
\newblock \emph{arXiv preprint arXiv:2012.07436}, 2021.
\newblock URL \url{https://arxiv.org/abs/2012.07436}.

\bibitem[Zivot \& Wang(2003)Zivot and Wang]{var}
Eric Zivot and Jiahui Wang.
\newblock Vector autoregressive models for multivariate time series.
\newblock \emph{Springer New York}, pp.\  385--429, 2003.
\newblock \doi{10.1007/978-0-387-32348-0_11}.
\newblock URL \url{https://doi.org/10.1007/978-0-387-32348-0_11}.

\end{thebibliography}

\documentclass{article} 

\usepackage{iclr2025_conference,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{graphicx}
\usepackage{times}
\usepackage{soul}
\usepackage{url}
\usepackage{tabularx}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{subfig,mwe}
\usepackage{algorithm}
\usepackage{bbm}
\usepackage{makecell}
\usepackage{times}
\usepackage{soul}
\usepackage{url}
\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{subfig,mwe}
\usepackage{graphicx}
\usepackage{subcaption} 
\usepackage{lipsum}
\usepackage{algorithm}
\usepackage{bbm}
\usepackage{makecell}
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{sidecap}
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{sidecap}
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\newtheorem{proposition}{Proposition}

\title{Efficient Linear Attention for Multivariate Time Series Modeling via Entropy Equality}

\author{
Mingtao Zhang$^{1}$, Guoli Yang$^{1}$, Zhanxing Zhu$^{2}$, Mengzhu Wang${^{1}}$, Xiaoying Bai${^{1}}$ \\
}

\begin{document}

\maketitle
\vspace{-5mm}
\begin{tabular}{c}
\textsuperscript{1} Advanced Institute of Big Data, Beijing\quad
\textsuperscript{2} University of Southampton\quad
\\
\vspace{0.5em}
\vspace{0.1cm}
\end{tabular} 

\begin{abstract}
Attention mechanisms have been extensively employed in various applications, including time series modeling, owing to their capacity to capture intricate dependencies; however, their utility is often constrained by quadratic computational complexity, which impedes scalability for long sequences. In this work, we propose a novel linear attention mechanism designed to overcome these limitations. Our approach is grounded in a theoretical demonstration that entropy, as a strictly concave function on the probability simplex, implies that distributions with aligned probability rankings and similar entropy values exhibit structural resemblance. Building on this insight, we develop an efficient approximation algorithm that computes the entropy of dot-product-derived distributions with only linear complexity, enabling the implementation of a linear attention mechanism based on entropy equality. Through rigorous analysis, we reveal that the effectiveness of attention in spatio-temporal time series modeling may not primarily stem from the non-linearity of softmax but rather from the attainment of a moderate and well-balanced weight distribution. Extensive experiments on four spatio-temporal datasets validate our method, demonstrating competitive or superior forecasting performance while achieving substantial reductions in both memory usage and computational time.
\end{abstract}

\section{Introduction}
Attention mechanisms have emerged as a pivotal component across a wide spectrum of modern machine learning applications~\cite{attention}, such as machine translation, image captioning~\cite{vit}, and time series forecasting~\cite{informer,Crossformer}. By enabling models to dynamically focus on relevant parts of the input, attention has significantly advanced the capacity to capture complex temporal dependencies in sequential data. Despite these successes, the widespread adoption of attention—especially in domains involving long sequences—is hindered by a critical limitation: its quadratic computational complexity with respect to sequence length~\cite{informer,linformer}. This scalability issue becomes particularly prohibitive in large-scale or real-time settings, including long-horizon time series forecasting and resource-constrained deployment environments, where both efficiency and accuracy are paramount.

Motivated by the pressing need to overcome the computational bottlenecks of conventional attention mechanisms, this work seeks to develop a scalable alternative that preserves its representational strength. As data volumes continue to grow and models are increasingly deployed in diverse and dynamic scenarios, the development of efficient attention variants has attracted growing interest. Drawing upon principles from information theory, we investigate how entropy—a measure of uncertainty and information content—can be leveraged to formulate a linear-time approximation of attention, enabling scalable modeling without compromising expressive power.

The main contributions of this paper are threefold. First, we establish a theoretical foundation by proving that entropy is a strictly concave function over the simplex of probability distributions. This property implies that distributions with similar entropy values exhibit structural similarity when their probability rankings align, thus offering a theoretical basis for approximating attention weights with reduced computational burden. Second, building on this insight, we propose an efficient approximation algorithm that computes the entropy of dot-product-induced distributions in linear time, dramatically lowering the computational overhead. Third, we introduce a practical linear attention module, termed Entropy-Aware Linear Attention (EALA), which incorporates entropy-based weighting to mimic the focus mechanism of standard attention. This module can be seamlessly integrated into existing attention-based architectures. Extensive experiments on four spatio-temporal datasets demonstrate that our approach achieves competitive—and at times superior—performance compared to state-of-the-art methods, while significantly reducing both memory consumption and inference time. Through these contributions, we aim to pave the way for more scalable and efficient attention-based models in resource-aware applications.

\section{Related Work}
\label{Related}
Traditional studies have primarily relied on the classical statistical models, \textit{e.g.} Autoregressive Integrated Moving Average (ARIMA)~\cite{ARMIA}, Vector Autoregressive (VAR)~\cite{var}, and Hidden Markov Models (HMMs). These models assume a strong linear dependency across timestamps. However, they fail to capture the complex patterns that require non-linear mapping. Recently, numerous deep learning-based methods have been proposed. Typical models can be categorized into RNN-based~\cite{lstm,gru} (which formulate time series as observations of generated by a sequence of latent states), TCN-like~\cite{tcn} (which use dilated CNNs to capture the temporal dependencies), Transformer-based~\cite{informer,Crossformer,STAEformer} (which leverages nonlocal attention weights to model temporal dynamics) and some other types, \textit{e.g.}, linear models~\cite{nbeats}. The above methods show different way for temporal dependency modeling, with RNN and TCN believed to be more efficient in capturing local pattern and Transformer structure superior in long-range dependency modeling.

Beyond the univariate forecasting, spatio-temporal time series forecasting shows a case where there are multiple time series connected as a graph. Graph neural networks (GNNs)~\cite{gcn} have been a popular choice in this scenario, enabling models to predict multivariate time series by jointly considering dependencies in temporal and spatial dimensions~\cite{STGCN,DCRNN,MTGNN}. Multiple models among them, such as STGCN~\cite{STGCN}, DCRNN~\cite{DCRNN}, \textit{etc.}, require a pre-defined graph representing the structural correlations between the series, which may be unavailable in real-world applications. Efforts to overcome this limitation include learning about the discrete graph structure~\cite{SGL,GTS} or weighted adjacency matrices~\cite{MTGNN,DGCRN,DFDGCN}. 
The learned graphs can be static~\cite{AGCRN,GWNet}, as matrices shared between different timestamps, or dynamic~\cite{DGCRN,dgcnn,MTGNN,PDformer}, which is commonly calculated as a similarity matrix based on separate inputs. Unlike the static graph learning which can be inflexible, dynamic graph learning is frequently used in recent research, with the adjacency matrices calculated as the similarity matrices. MTGNN~\cite{MTGNN} designs different types of learned graph, in which the dynamic graph is constructed by dot product on the series embedding. D2STGNN~\cite{D2STGNN} combines the pre-defined graph matrix with the attention matrices to model the dynamic relationship.

Therefore, attention~\cite{attention} has been utilized in various applications, both for temporal modeling and spatial correlation modeling. However, the quadric complexity has always been a harmful issue for attention in scalable forecasting. Here we give an efficient implementation of attention.

\section{Methodology}
\label{Methodology}

Firstly, we formalize the importance of entropy-equality, which means that when the ranking of the probability is the same, similar entropy will lead to similar distribution. Secondly, we propose an efficient approximating algorithm, which needs only linear complexity to calculate the entropy of the distribution derived from dot-production. Finally, we propose an implementation of linear attention based on the entropy-equality.

Entropy acts as a global metric of probability distribution, and similar distribution will have similar entropy. Therefore, approximating a distribution intuitively requires the alignment of entropy. Moreover, on the opposite direction, it can be proved that under certain circumstances, equality of entropy can result in the similar distribution, as following formalized.
\begin{proposition}
Let $\Delta^n$ denote the $n$-dimensional probability simplex defined as:
\[
\Delta^n = \left\{ \mathbf{p} = (p_1, p_2, \ldots, p_n) \in \mathbb{R}^n : p_i \geq 0 \ \forall i, \ \sum_{i=1}^n p_i = 1 \right\}.
\]
The entropy function $H: \Delta^n \to \mathbb{R}$ is given by:
\[
H(\mathbf{p}) = -\sum_{i=1}^n p_i \log p_i,
\]
where $\log$ denotes the natural logarithm. And the KL-divergence is defined as:
\[
KL(\mathbf{q}||\mathbf{p}) = \sum_{i=1}^n q_i \log \frac{q_i}{p_i},
\]
Then the following properties hold:
\textbf{Implication under Consistent Ordering:} If two probability distributions $\mathbf{p}, \mathbf{q} \in \Delta^n$ have the same ordering (i.e., for all indices $i,j$, $p_i \leq p_j$ if and only if $q_i \leq q_j$), then:

\begin{align*}
    KL(\mathbf{q}||\mathbf{p}) &= (H(\mathbf{q})-H(\mathbf{p}))+\sum_i(p_i-q_i)\log p_i\\
 &\le |H(\mathbf{q})-H(\mathbf{p})|+|\sum_i(p_i-q_i)\log p_i|
\end{align*}

Here $p_i$ is the prior distribution, which is fixed and we assume it to be non-zero. Then $|\sum_i(p_i-q_i)\log p_i|$ is a linear function for $q_i$ and reaches its maximum on border of $\Delta^n$. As they have the same ranking, we therefore point out that KL-divergence between $\mathbf{p}$ and $\mathbf{q}$ can be bounded and mostly the maximum cannot be reached (as the border of $\Delta^n$ means degenerate one-hot distribution).

\end{proposition}

Therefore, we first loosen the requirement of $p_i\ge 0$ as it is generally satisfied, and point out that, when a distribution is based on scores $\mathbf{x}$ as $\phi(\mathbf{x})$, and the distribution is positively correlated, we can approximate a difficult $\phi'(\mathbf{x})$ by tuning $\theta$ parameter in $\phi_\theta(\mathbf{x})$, to make their entropy equal ($\phi',\phi_\theta$ are both monotonic). Even if $\phi_\theta$ is not a strictly distribution, this is also a reasonable heuristic approximation.

Therefore, we secondly propose an efficient approximation of entropy, when the distribution is derived from dot-production. We take the notation from attention and omit the commonly utilized scaling factor, formalizing attention as $Attn(q,k_j)=softmax(q\cdot k_j)=\frac{\exp(q\cdot k_j)}{\sum_l\exp(q\cdot k_l)}$, which forms a distribution for each $q$ across different $k_j$. The entropy is therefore calculated as:
\begin{equation}\label{entropy1}
    \begin{aligned}
    H(Attn(q,\cdot))&=-\sum_j Attn(q,k_j)\log(Attn(q,k_j))\\
    &=-\sum \frac{\exp(q\cdot k_j)}{\sum_l\exp(q\cdot k_l)}(q\cdot k_j-\log\sum_l\exp(q\cdot k_l))\\
    &=-\sum \frac{\exp(q\cdot k_j)q\cdot k_j}{\sum_l\exp(q\cdot k_l)}-\log\sum_l\exp(q\cdot k_l)
\end{aligned}
\end{equation}

As our goal is to approximate the result, we utilize first-order Taylor expansion for efficient calculation. 
After the expansion, the result is approximated by:

\begin{equation}\label{entropy2}
    H(Attn(q,\cdot))\approx-\sum_j \frac{(1+q\cdot k_j)q\cdot k_j}{N+\sum_lq\cdot k_l}-\log(N+\sum_lq\cdot k_l)
\end{equation}

As $\sum_jq\cdot k_j=q\sum_jk_j$, it can be easily calculated for all $q$ with linear complexity, avoiding the requirement to calculating all $q\cdot k_j$. However, it still faces the challenging calculation of $\sum_j(q\cdot k_j)^2$. We re-formalize it to show it can also be calculated with linear complexity.

\begin{align*}
    \sum_j(q\cdot k_j)^2&=\sum_j qk_j^Tqk_j^T=\sum_j qk_j^T(qk_j^T)^T\\
    &=\sum_j qk_j^Tk_jq^T=q(\sum_j k_j^Tk_j)q^T
\end{align*}

As $(\sum_j k_j^Tk_j)$ can also be calculated without separately calculating each $q\cdot k_j$, this calculation can be further re-formalized as matrix multiplication form. By replacing $\sum_j(q\cdot k_j)$ and $\sum_j(q\cdot k_j)^2$ in the entropy, we achieve the efficient calculation of entropy without separately calculating each $q\cdot k_j$. This calculation trick can be utilized in multiple fields, and here we propose to implement an efficient linear attention mechanism based on the proposition.

To approximate distribution $\phi'(\mathbf{x})=softmax(\mathbf{x}),x_j=Attn(q,k_j)$, we should design a new distribution $\phi_\theta(\mathbf{x})$ with tunable parameter $\theta$. Moreover, $\phi_\theta$ should also be positively correlated with $\mathbf{x}$. We first design a simple linear function $\phi_\theta^0(\mathbf{x})=1+\frac{\mathbf{x}}{\theta_q}$, and $\phi_\theta(\mathbf{x})=Norm(\phi_\theta^0(\mathbf{x}))$ is normalized to make their summarization as $1$. Therefore, we approximate $H(\phi')=H(Attn)$ by $H(\phi_\theta)$. Meanwhile, as we used first-order Taylor expansion, we try to minimize the errors, and therefore, we should turn the meaning of $q\cdot k_j$ as $0$, implemented by $\hat{k}_j=k_j-\bar{k}_j$, with $\bar{k}_j$ as the average of each $k_j$. Therefore, in implementation, $\theta$ is searched to satisfy the following equation:

\begin{align}\label{entropy3}
    H(\phi_\theta)=H(\phi') \Rightarrow \forall q, &H(Norm(1+\frac{q\cdot \hat{k}_j}{\theta_q}))=H(Attn(q,\cdot))
\end{align}

Finally the optimal $\theta$ has analytic solution as $\theta_q^*=\sqrt{\frac{\sum_j (q\cdot k_j)^2}{2*N*(\log N+H(Attn(q,c\dot)))}}+\epsilon$, which is also linearly computable, with $\epsilon$ to avoid division by $0$ in $1+\frac{\mathbf{x}}{\theta^*}$. The detailed calculation and proof in Appendix.

Therefore, our model can be generally formalized as in Algorithm~\ref{attention}.

\begin{algorithm}
\caption{Entropy-Equal Linear Attention}\label{attention}
\begin{algorithmic}[1]
\Require Input $Q=[q_1,\cdots,q_N],K=[k_1,\cdots,k_N],V=[v_1,\cdots,v_N],q_i,k_i,v_i\in \mathbb{R}^C$
\Ensure Output $O=[o_1,\cdots,o_N]$

\State Calculate $H(Attn(q_i,\cdot))$ by Equation~\ref{entropy1} and~\ref{entropy2}
\State $\forall q_i$, calculate $\theta^*_{q_i}=\sqrt{\frac{\sum_j (q\cdot k_j)^2}{2*N*(\log N+H(Attn(q_i,\cdot)))}}+\epsilon$ 
\If{$C>N$}
    \State $o_i \gets (\sum_j(1+\frac{q_i\cdot k_j}{\theta^*_{q_i}})v_j)/N$ \Comment{For small scale dataset, with complexity $O(N^2C)$}
\Else
    \State $o_i \gets (\sum_jv_j+q_i/\theta^*_{q_i}\cdot(k_j\cdot v_j))/N$ \Comment{For large scale dataset, with complexity $O(NC^2)$}
\EndIf
\State \Return $O=[o_1,\cdots,o_N]$
\end{algorithmic}
\end{algorithm}

Based on our proposed linear attention mechanism, we propose an attention-only models for spatio-temporal forecasting, named as ELinFormer (Entropy-Equal Linear Transformer), and the base version is attention-only model without linear implementation. 

\section{Experiments}
\label{Experiments}

\subsection{Experimental Settings}

\subsubsection{Datasets} We evaluated performance on real-world public datasets, including five highway traffic datasets (PEMS-BAY, PEMS03, PEMS04, PEMS07, PEMS08~\cite{pems}). The PEMS datasets are all collected from the Caltrans Performance Measurement System, and they represent the cases where there are intuitively strong correlations between series, so the spatial and temporal dependency modeling are both vital. Detailed information is provided in Table~\ref{Datasets}.

\begin{table}[h!]
    \centering
    
    \caption{Data Description}\label{Datasets}
    \resizebox{\linewidth}{!}{
    \begin{tabular}{cccccc}
    \toprule
        \textbf{Datasets} & \textbf{\#Nodes} & \textbf{\#Edges} & \textbf{\#Timesteps} & \textbf{\#Time Interval} & \textbf{Time range} \\ \midrule
        \textbf{PEMS-BAY} & 325 & 2369 & 52128 & 5min & 01/01/2018-06/30/2018   \\ 
        \textbf{PEMS03} & 358 & - & 26208 & 5min & 05/01/2012-07/31/2012   \\
        \textbf{PEMS04} & 307 & 340 & 16992 & 5min & 01/01/2018-02/28/2018   \\ 
        \textbf{PEMS07} & 883 & 866 & 28224 & 5min & 05/01/2017-08/31/2017   \\ 
        \textbf{PEMS08} & 170 & 295 & 17856 & 5min & 07/01/2016-08/31/2016  \\
         \bottomrule
    \end{tabular}
}
    \label{Dataset}
\end{table}

\begin{table}[htb]
\caption{Comparison of forecasting results of the Vallina and our attention on the datasets. The results are the best recording collected from BasicTS or the original paper, with the other unavailable results tested in our setting.\protect\footnotemark[1]}
\label{Compare result}
\resizebox{\textwidth}{!}{
\begin{tabular}{c|c|ccc|cc}
\toprule
       Dataset & Model& MAE         & RMSE              & MAPE      &Memory (MiB) &Time (s/epoch)   \\
                             \midrule
\multirow{3}{*}{PEMS03} &STAEformer(Baseline) &15.35 & 27.55 & \textbf{15.18\%}& 8336& 186  \\
& ELinFormer(Ours)&\textbf{ 14.81}&\textbf{ 24.86} & 15.21\% &5278 &151 \\
& Improvement $\Delta$ & 3.5 & 9.8&-0.2 & 36.7\%& 18.8\% \\
\cmidrule(lr){1-7}
\multirow{3}{*}{PEMS04} &STAEformer(Baseline) &\textbf{18.22} & 30.18 & 11.98\%& 7096& 101  \\
& ELinFormer(Ours)& \textbf{18.22}& \textbf{29.82} & \textbf{11.92\%} &5098 &86 \\
& Improvement $\Delta$ & 0.0 & 1.2&0.5 & 28.2\%& 14.9\%\\
\cmidrule(lr){1-7}
\multirow{3}{*}{PEMS07} &STAEformer(Baseline) &19.14 & 32.60 & \textbf{8.01}\%& >10240& -  \\
& ELinFormer(Ours)& \textbf{19.06}& \textbf{32.39} & 8.04\% &9948 &375 \\
& Improvement $\Delta$ & 0.4 & 0.6&-0.4 & -& - \\
\cmidrule(lr){1-7}
\multirow{3}{*}{PEMS08} &STAEformer(Baseline) &\textbf{13.46} & \textbf{23.25} & \textbf{8.88}\%& 4470& 59  \\
& ELinFormer(Ours)& 13.51& 23.37 & 8.90\% &3834 &55 \\
& Improvement $\Delta$ & -0.4 & -0.9& -0.2 & 14.2\%& 6.8\% \\
\bottomrule
\end{tabular}
}
\end{table}

\begin{table}[htb]

\caption{Overall forecasting results of the models on the highway traffic datasets. `\_' is used to highlight the second best-performing model. The results are the best recording collected from BasicTS or the original paper, with the other unavailable results tested in our setting.\protect\footnotemark[1]}
\label{Main result}
\resizebox{\textwidth}{!}{
\begin{tabular}{l|ll|ll|ll|ll|ll}
\toprule
       \multirow{2}{*}{Model} & \multicolumn{2}{c|}{PEMS-BAY} & \multicolumn{2}{c|}{PEMS03} & \multicolumn{2}{c|}{PEMS04} & \multicolumn{2}{c|}{PEMS07}& \multicolumn{2}{c}{PEMS08} \\
&MAE&RMSE&MAE&RMSE&MAE&RMSE&MAE&RMSE&MAE&RMSE\\
\midrule
AGCRN&1.60&3.67&15.24&26.65&19.28&31.02&20.68&34.45&15.78&24.76\\
DCRNN&1.59&3.69&15.54&27.18&19.66&31.18&21.16&34.15&15.23&24.17\\
GTS&1.68&3.79&15.41&26.15&21.32&33.55&22.47&35.42&16.92&26.68\\
D2STGNN&\textbf{1.52}&\underline{3.53}&\underline{14.63}&26.31&18.32&29.89&19.49&32.59&14.10&23.36\\
HimNet&-&-&15.11&26.56&\textbf{18.14}&\underline{29.88}&19.21&32.75&13.57&\textbf{23.22}\\
\cmidrule(lr){1-11}
STGCN&1.63&3.72&15.83&27.51&19.76&31.51&22.25&35.83&16.19&25.51\\
MTGNN&1.60&3.71&14.85&\underline{25.23}&19.13&31.03&21.01&34.14&15.25&24.22\\
GWNet&1.59&3.68&\textbf{14.59}&25.24&18.53&29.92&20.44&33.38&14.67&23.55\\
\cmidrule(lr){1-11}

PDFormer&1.57&3.58&14.94&25.39&18.36&30.03&19.83&32.87&13.58&23.41\\
STWave&-&-&14.93&26.50&18.50&30.39&19.94&33.88&\textbf{13.42}&23.40\\
STAEformer&1.54&\textbf{3.52}&15.35&27.55&\underline{18.22}&30.18&\underline{19.14}&\underline{32.60}&\underline{13.46}&\underline{23.25}\\
ELinFormer&1.53&\textbf{3.52}&14.81&\textbf{24.86}&\underline{18.22}&\textbf{29.82}&\textbf{19.06}&\textbf{32.39}&13.51&23.37\\
\bottomrule
\end{tabular}
}
\end{table}

\subsubsection{Baselines} We identify STAEformer~\cite{STAEformer} as the base version of attention-only models, which is the state-of-the-art in current spatio-temporal forecasting. For a complete comparison, we select multiple types of baseline models. We exclude some baselines like STDMAE~\cite{STD-MAE}, STEP~\cite{STEP} (as they utilize pretraining strategy), STDN~\cite{STDN} (as it focuses on unusual setting of decomposition) and BigST~\cite{BigST} (as it focuses on extremely small models) for fair comparison. 

(1) \textit{Models with RNN-like Backbones}: DCRNN ~\cite{DCRNN}, AGCRN~\cite{AGCRN}, GTS~\cite{GTS}, D2STGNN~\cite{D2STGNN} and HimNet~\cite{HimNet} are chosen as baselines of the RNN-like backbones. 

(2) \textit{Models with TCN-like Backbones}: We choose STGCN~\cite{STGCN}, Graph Wavenet (abbreviated as \textit{GWNet})~\cite{GWNet}, and MTGNN~\cite{MTGNN}, as representatives of models that incorporate TCN-like backbones. 

(3) \textit{Models with Transformer-like Backbones}: PDFormer \cite{PDformer} and STAEformer~\cite{STAEformer} are chosen as representatives that incorporate Transformer-like backbones. Since the algorithm generating the mask of PDFormer is unavailable, the missing results will be omitted in the table and analysis.

(4) \textit{Other Attention Methods} Moreover, we compare our module with other implementation of attention-like module: Vallina-Attention~\cite{attention}, and previous Linformer~\cite{linformer}, on various backbones to specifically assess the impact of dynamic graphs. We exclude the comparison with DGCRN~\cite{DGCRN}, D2STGNN~\cite{D2STGNN}, because their graphs are specific forms of the self-attention mechanism, and their temporal parts do not include attention.

\subsubsection{Implementation Details} We utilize the pre-processing pipeline of the original method settings. Specifically, we use the framework of BasicTS~\cite{exploring}, and conduct experiments on an Intel(R) Xeon(R) Platinum 8255C CPU@2.50GHz and an NVIDIA RTX 3080 GPU, using PyTorch version 1.10.0, Python 3.8 and CUDA 11.3. The learning rate, optimizer, and batch size are set according to the backbone configurations. For datasets without an optimal configuration, we select the most widely used. We assign $\epsilon$ as $1e-8$ to align with other baselines' configuration. But we suggest that utilizing large $\epsilon$ may stabilize the training process, as inspired by the empirical trick in Large Language Model training.

\subsubsection{Evaluation Metrics} To ensure consistency with previous studies, we utilize two metrics in our experiments: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE) and Mean Absolute Percentage Error (MAPE). Missing values are excluded from the calculation of these metrics.

\subsection{Main Results}

We first conducted a further comparison between the backbones equipped with other dynamic graphs and with our dynamic graphs to demonstrate the performance of various graph learning modules on different backbones. As presented in Table~\ref{Compare result}, the following observations can be inferred:

(1) Our linear implementation of attention achieves comparable or even better results, with significantly reduced memory and time complexity. Notably we got better results for some models in PEMS03 compared to original papers: STAEformer got $14.94$(MAE) and $25.64$(RMSE). We assume it may due to the differences of data distribution from validation set, making the choice of best model erroneous. In our implementation, the improvement of our model over the Vallina attention in STAEformer in PEMS03 is reduced to $0.9\%$ and $3.0\%$. This result is more reasonable because our model tries to approximate the distribution of Vallina attention. Minor improvements is possible for more robust implementation with possibly faster training, but significant improvement in performance is not reasonable.

(2) When the dataset becomes larger, the efficiency improvements in memory and time is more significant. It is because the FFN Layers in attention become prominent when the dataset becomes smaller but the feature dimension is fixed. Besides, our model mostly improves the efficiency in spatial dependency modeling, because the look-back window in spatio-temporal forecasting is often small, even less than feature dimension in attention. Under this circumstances, our model does minor improvement for temporal dependency modeling. 

(3) The experiment results and the above analysis can demonstrate the general role of attention in spatio-temporal modeling. From the performance analysis, attention may not improve the model by its non-linearity, because an approximating linear function can almost achieve the same result. Instead, we believe attention makes its contribution based on the more expressive distribution, which adaptively choose the suitable level of sharpness for the weighted average afterwards. Besides, the efficiency  

We then compared the linear methods to popular baselines in spatio-temporal forecasting, as presented in Table~\ref{Main result}. It should be stressed that our aim is not to propose a perfect forecasting model, but to show an efficient substitute to the current attention, and therefore analyze the effectiveness of attention in spatio-temporal forecasting. Based on the results, the following observations can be made: 

(1) Transformer often gets the optimal results in various datasets, probably due to its dynamic modeling capability. The performance of backbones that capture only static dependencies, e.g. RNN that only captures the dependencies between the previous timestamp with the current one, get inferior performance.

(2) Among the above methods, out model gets comparable or even better results compared to the current state-of-the-art. What's more, our model provides a unified framework for multiple datasets, with minimal heuristic designs and external information. Our model does not require predefined graphs or specifically designed threshold and mask, and is easy to transfer to more nodes and other datasets with less hyperparameter tuning.

(3) Our model provides an efficient implementation for current spatio-temporal forecasting. Most current methods rely on GNN or spatial attention to model the spatial correlation, which require $O(N^2d)$ complexity which may be hard to scale. PDFormer~\cite{PDformer} and Sparse Graph Learning~\cite{SGL} propose more efficient strategy with ideal complexity $O(|E|)$ with $|E|$ as the edge number in the graph, but it may be challenging for CUDA implementation and speeding, and the upperbound of complexity can be $O(N^2)$. STWave~\cite{STWave} utilizes frequency filtering before attention, with the complexity $O(NT^2+TN\log N)$, but it still has a super-linear complexity. 

Therefore, our model achieves impressive performance, with linear complexity. It utilizes only simple multiplication and addition, and is therefore possibly to utilize CUDA kernel optimization for acceleration, e.g. utilizing \textit{torch.baddbmm} function to incorporate both multiplication and addition into less operations. However, as this is outside the scope of our research, so we leave this issue for future investigation. 

\section{Conclusion}

We propose a simple implementation of linear attention, which improves spatio-temporal forecasting by linearly approximates the balanced weight distribution from the Vallina attention. It replaces spatial and temporal attention in forecasting by a linear counterpart. In the process, we analyze the relationship between entropy and the global distribution, to show both intuitively and theoretically that entropy-equality leads to similar distribution. We further propose a novel efficient algorithm to linearly approximate the entropy derived from the softmax after matrix multiplication. We therefore propose a linear implementation of attention based on the above conclusion.
Moreover, we integrate the module into various the current attention-only model. Experimental results on benchmark datasets demonstrate the module's effectiveness in improving performance and efficiency. Furthermore, the results highlight the potential for analyzing the role of attention as non-linear transformation and efficient weighted average, which can be further investigated. In the future, we shall test the linear attention in long-range time series modeling and other aspects like natural language processing (NLP).



\bibliography{iclr2025_conference}
\bibliographystyle{iclr2025_conference}

\appendix
\section{Appendix}
You may include other additional sections here.

\end{document}